# Gaussian Process Priors for Boundary Value Problems of Linear Partial Differential Equations

### v1

@@ 3.3 B-EPGP with linear boundaries
We consider the 2-D wave equation with Dirichlet boundary condition at $x=0$ and Neumann boundary condition at $y=0$:
$\begin{cases}\Box u(x,y,t)=0&(x,y)\in(0,\infty)^{2},t>0\\ u(0,y,t)=0&y\in(0,\infty),t>0\\ u_{y}(x,0,t)=0&x\in(0,\infty),t>0\end{cases}$
The EPGP basis in full space for $\square u=0$ is given by
$\{e^{\alpha x+\beta y+\tau t}\colon\alpha,\beta,\tau\in\mathbb{C},\,\alpha^{2}+\beta^{2}=\tau^{2}\}.$
Performing the algorithm from the beginning of the section leads to the basis:
@@
which satisfies both boundary conditions at once. Both the calculation of the basis and an implementation are described in Appendix G.2. ∎
@@
Consider the 1-D wave equation on a line segment $(0,\pi)$ with Dirichlet boundary at both $x=0$ and $x=\pi$,
@@
An EPGP basis for the equation in full space is $e^{\sqrt{-1}\xi(x\pm t)}$ for $\xi\in\operatorname*{\mathbb{R}}$. To construct a basis that also adheres to the boundary condition at $x=0$, we use Section 3.2 to obtain $e^{\pm\sqrt{-1}\xi t}\sin(\xi x).$ Choosing $\xi\in\mathbb{Z}$ gives a basis which satisfies the condition at $x=\pi$ as well. Both the calculation of the basis and an implementation are described in Appendix G.1. ∎
Concerning Example 3.7, we have:
@@
B-EPGP gives $\{e^{\sqrt{-1}jx}\sin(jx)\}_{j\in\mathbb{Z}}$ as a basis for (4). Its span is dense in the solution space of (4).
@@
Consider the equation in Example 3.6 in a square with Dirichlet boundary conditions:
$\begin{cases}\square u=0&\text{for }x,y\in(0,\pi),\,t>0\\ u=0&\text{if $x$ or $y=0$ or $\pi$.}\end{cases}$ (5)
@@
In this case we retrieve the method of separation of variables (Fourier series approach). The calculation of the basis and an implementation are described in Appendix F.1. ∎
@@
B-EPGP gives the basis (6) for (5). The span of (6) is dense in the solution space of (5).
In Theorem B.1 we will give a general statement which incorporates both Theorems 3.8 and 3.10.
@@ 3.4 Hybrid B-EPGP
In its current form, B-EPGP can be implemented with varying degrees of effort for many linear boundary conditions. For purely curved surfaces, we can use the direct method in Section 3.1 to obtain less accurate results. For domains which have both curved and flat pieces of the boundary, we can use B-EPGP to fulfill the linear boundary conditions and sample data at the curved pieces.
Example 3.11 (Sector).
@@
Consider the space-time domain $\Omega=\{(x,y,t)\colon x^{2}+y^{2}<1,x>0,y>0,t>0\}$ which has the spatial boundary $\Gamma=\left([0,1]\times\{0\}\right)\cup\left(\{0\}\times[0,1]\right)\cup\{(\cos\theta,\sin\theta)\colon\theta\in[0,\pi/2]\}$, represented in Figure 2.
$x$$y$$\Gamma$
Figure 2: Circular sector in the two space directions of $\Omega$ delimited by the boundary $\Gamma$, as used in Example 3.11.
We consider the 2-D wave equation in this circular sector
$\begin{cases}\square u=0&\text{in }\Omega\\ u=0&\text{on }\Gamma.\end{cases}$
To account for the flat pieces of $\Gamma$, B-EPGP gives a basis similar to the one in Example 3.6, namely
$e^{\alpha x+\beta y+\tau t}-e^{-\alpha x+\beta y+\tau t}-e^{\alpha x-\beta y+\tau t}+e^{-\alpha x-\beta y+\tau t},$
@@
where $\alpha,\beta,\tau\in\mathbb{C}$ and $\alpha^{2}+\beta^{2}=\tau^{2}$. To account for the circular piece, we sample many points $(\cos\theta,\sin\theta,t)$ and restrict our Gaussian process to satisfy $u=0$ at those points. An implementation can be found in Appendix 5.3. ∎
4 Experimental Comparison to the State of the Art
We want to compare with the Neural Operators from (Raonic et al., 2024), namely CNO and FNO. The specific experiment setting we use is 1-D wave equation with Neumann boundary condition:
$\begin{cases}u_{tt}=u_{xx}&\text{in $(0,\infty)\times(0,\infty)$}\\ u(0,x)=h_{1}(x)&\text{for $x\in[0,\infty)$}\\ \frac{\partial}{\partial t}u(0,x)=h_{2}(x)&\text{for $x\in[0,\infty)$}\\ \frac{\partial}{\partial x}u(t,0)=0&\text{for $t\in(0,\infty)$}\end{cases}$
where $h_{1}(x)=f(x-3)+f(x+3)+g(x-1)+g(x+1)$ and $h_{2}(x)=f^{\prime}(x-3)-f^{\prime}(x+3)$ for $f(x)=e^{-5x^{2}}$, $g(x)=e^{-10x^{2}}$. The exact solution is given by
$u(x,t)$ $=f(x+t-3)+f(x-t+3)$
$\ +\frac{1}{2}g(x+t-1)+\frac{1}{2}g(x-t-1)$
$\ +\frac{1}{2}g(x+t+1)+\frac{1}{2}g(x-t+1)$
$=e^{-5(x+t-3)^{2}}+e^{-5(x-t+3)^{2}}$
$\ +\frac{1}{2}e^{-10(x+t-1)^{2}}+\frac{1}{2}e^{-10(x-t-1)^{2}}$
$\ +\frac{1}{2}e^{-10(x+t+1)^{2}}+\frac{1}{2}e^{-10(x-t+1)^{2}}.$
For B-EPGP, we can easily implement the Neumann boundary into our basis, which means our basis is:
$\{e^{\alpha(x\pm t)}+e^{\alpha(-x\pm t)}\}_{\alpha\in\mathbb{C}}.$
See Appendix A for details on how to calculate the basis. We consider both having $n=121$ and $n=1201$ sample points from initial condition on the interval $[0,12]$, and for EPGP, CNO and FNO, we sample a point on the boundary at the interval of $0.2$. We show that B-EPGP is between one and two orders of maagnitude superior to the neural operator methods in Table 1.
Table 1: This table shows the median pointwise absolute and relative $L_{2}$ error in $[0,4]\times[0,8]$ for the experimental setting described in Section 4.
@@ 4 Experimental Comparison to the State of the Art
(a) EPGP
(b) B-EPGP (ours)
(c) CNO
(d) FNO
(e) EPGP Error
(f) B-EPGP (ours) Error
(g) CNO Error
(h) FNO Error
Figure 3: Estimated solution of 1-D wave equation with Neumann boundary condition on the top part of the domain. This condition represents the reflection of the wave by the boundary. We use $n=1201$ and four different algorithms out of which ours perform better by at least one order of magnitude. We stress that the error in the neural operator approaches increases over time, whereas the error in B-EPGP stays almost constant.
5 Examples
In this section we report the results of some of our most relevant examples outlined theoretically in Section 3. Results for the other examples in Section 3 and more can be found in Appendices F.1F.2F.3G.1G.3HI.
There are two main ways in which we check the accuracy of our numerical solutions:
  • Verify our solvers in contexts when we know the unique true solutions to the PDEs.
  • Check the conservation of energy for the wave equation in bounded domains (see Appendix E).
The latter is an important physical property which guarantees that both the equation and the (Dirichlet or Neumann) boundary conditions are satisfied.
In Section 4 we saw that our old algorithm EPGP and our new algorithm B-EPGP perform much better than the state of the art. In the following examples, we will compare EPGP and B-EPGP directly.
5.1 Comparison between B-EPGP and EPGP
Consider the following initial boundary value problem for the 2-D wave equation in a halfspace $x>0$:
$\begin{cases}u_{tt}=u_{xx}+u_{yy}&\text{for $t\in(0,\infty),x\in(0,\infty),y\in\operatorname*{\mathbb{R}}$}\\ u(0,x,y)=f(x,y)&\text{for $x\in[0,\infty),y\in\operatorname*{\mathbb{R}}$}\\ \frac{\partial}{\partial t}u(0,x,y)=0&\text{for $x\in[0,\infty),y\in\operatorname*{\mathbb{R}}$}\\ \frac{\partial}{\partial x}u(t,0,y)=0&\text{for $t\in(0,\infty),y\in\operatorname*{\mathbb{R}}$}\end{cases}$
with initial condition $f=f_{1}+f_{2}+f_{3}$ for
$f_{1}(x,y)=$ $J_{0}(5\sqrt{((x-1)^{2}+(y-1)^{2})})$
$+J_{0}(5\sqrt{((x+1)^{2}+(y-1)^{2})}),$
$f_{2}(x,y)=$ $J_{0}(10\sqrt{((x-2)^{2}+(y-2)^{2})})$
$+J_{0}(10\sqrt{((x+2)^{2}+(y-2)^{2})}),$
$f_{3}(x,y)=$ $J_{0}(5\sqrt{((x-3)^{2}+(y-3)^{2})})$
$+J_{0}(5\sqrt{((x+3)^{2}+(y-3)^{2})}),$
where $J_{0}$ is the Bessel function of order 00.
For this problem, the unique exact solution is given by
$u(t,x,y)$ $=f_{1}(x,y)\cos(5t)+f_{2}(x,y)\cos(10t)$
$+f_{3}(x,y)\cos(5t)$
We can use B-EPGP to obtain a basis which satisfies both the Neumann boundary condition at $x=0$ but also the initial speed at $t=0$, cf. Example 3.6:
$e^{\alpha x+\beta y+\tau t}+e^{-\alpha x+\beta y+\tau t}+e^{\alpha x+\beta y-\tau t}+e^{-\alpha x+\beta y-\tau t},$
where $\alpha,\beta,\tau\in\mathbb{C}$ and $\alpha^{2}+\beta^{2}=\tau^{2}$.
Table 2: The following results from Section 5.1 show the superiority of our approach B-EPGP over EPGP when comparing both the median L1-difference to the exact solution and the computation time. We perform ten repetitions and report the standard deviation.
Algorithm Abs Err($10^{-4}$) Rel Err(%) Time(s)
EPGP $4.26\pm 0.21$ $1.82\pm 0.09$ $6059$
B-EPGP (ours) $\textbf{0.48}\pm 0.02$ $\textbf{0.72}\pm 0.03$ $804$
We fix the number of basis elements to be the same. Hence, the number of parameters needed to be trained in the B-EPGP approach is reduced by a factor of 4 times over those of EPGP. Furthermore, the data on the boundary in EPGP increases the covariance matrix by a factor of 9, since both initial data, initial speed and boundary data need is necesarry. This triggers a significant improvement on both the running time and memory requirement. The results of B-EPGP in comparison to EPGP are shown in Table 2. A one-sided Paired Wilcoxon Signed Rank Test, with alternative hypothesis: true location shift is less than 0, shows that B-EPGP has a highly significant improvement with $p$-value of $0.0009766\ll 0.05$. In this experiment, B-EPGP reduces both running times and memory cost by one order of magnitude, while also improving the quality of the results by one order of magnitude.
(a) $t=0.0$,
EPGP
(b) $t=0.5$,
EPGP
(c) $t=1.0$,
EPGP
(d) $t=1.5$,
EPGP
(e) $t=2.0$,
EPGP
(f) $t=2.5$,
EPGP
(g) $t=3.0$,
EPGP
(h) $t=3.5$,
EPGP
(i) $t=4.0$,
EPGP
(j) $t=0.0,$
B-EPGP
(k) $t=0.5,$
B-EPGP
(l) $t=1.0,$
B-EPGP
(m) $t=1.5,$
B-EPGP
(n) $t=2.0,$
B-EPGP
(o) $t=2.5,$
B-EPGP
(p) $t=3.0,$
B-EPGP
(q) $t=3.5,$
B-EPGP
(r) $t=4.0,$
B-EPGP
Figure 4: Predictions to the 2-D wave equation with Neumann boundary on a halfspace made using EPGP and B-EPGP. We use this experiment as a benchmark as we can compare both predictions with a highly nontrivial exact solution. Table 2 shows that B-EPGP is better by an order of magnitude in both accuracy and computational time. Animations can be found in the ancillary files 2DBessel-S.mp4 and 2DBessel-R.mp4.
5.2 Wave in 3D
The 3-D wave equation is computationally challenging even without boundary conditions (see Appendix D which is concerned with adjusting EPGP to fit both initial condition and speed and comparison with the literature).
We use B-EPGP for 3-D wave equation with Neumann boundary condition on the planes $y=0$ and $z=0$ and initial conditions as follows.
$\begin{cases}u_{tt}=u_{xx}+u_{yy}+u_{zz}&\text{ for }x\in\operatorname*{\mathbb{R}},y,z,t>0\\ u=e^{-5r_{1}^{2}}+e^{-5r_{2}^{2}}+e^{-5r_{3}^{2}}&\text{at }t=0\\ u_{t}=0&\text{at }t=0\\ u_{y}=0&\text{at }y=0\\ u_{z}=0&\text{at }z=0\end{cases}$
where $r_{1},r_{2},r_{3}$ denote Euclidean distances between $(x,y,z)$ and $(1,1,1),(1,-1,1),(1,1,-1)$, respectively.
The B-EPGP basis we use is
$\ e^{ax+by+cz+dt}+e^{ax-by+cz+dt}+e^{ax+by-cz+dt}$
$+e^{ax-by-cz+dt}+e^{ax+by+cz-dt}+e^{ax-by+cz-dt}$
$+e^{ax+by-cz-dt}+e^{ax-by-cz-dt}$
which can be computed using the ideas in Section G.2.
There are not many other papers which tackle this and we suspect that computation of solutions for PDE in 4-dimensions incurs a curse of dimensionality. For instance, EPGP goes out of memory on an Nvidia A100 80GB. Other related papers either do not tackle wave equations with boundary conditions (Henderson et al., 2023) or do not tackle PDEs in 4-D (Raonic et al., 2024). In turn, our B-EPGP finishes with a low L1-error of 0.00088.

### v2

@@ 3.2 B-EPGP with polygonal boundaries
We consider the 2D wave equation with Dirichlet boundary condition at $x=0$ and Neumann boundary condition at $y=0$:
$\begin{cases}u_{tt}-u_{xx}-u_{yy}=0&(x,y)\in(0,\infty)^{2},t>0\\ u(0,y,t)=0&y\in(0,\infty),t>0\\ u_{y}(x,0,t)=0&x\in(0,\infty),t>0\end{cases}$
The EPGP basis for $u_{tt}-u_{xx}-u_{yy}=0$ is given by $\{e^{\alpha x+\beta y+\tau t}\colon\alpha,\beta,\tau\in\mathbb{C},\,\alpha^{2}+\beta^{2}=\tau^{2}\}.$ Performing the algorithm from the beginning of the section leads to the basis:
@@
which satisfies both boundary conditions at once. The calculation of the basis and an implementation with a $45^{\circ}$ wedge are described in Appendices J.2J.3, respectively. ∎
@@
Consider the 1D wave equation on a line segment $(0,\pi)$ with Dirichlet boundary at both $x=0$ and $x=\pi$,
@@
An EPGP basis for this equation is $e^{\sqrt{-1}\xi(x\pm t)}$ for $\xi\in\operatorname*{\mathbb{R}}$. Section 3.1 yields the basis $e^{\pm\sqrt{-1}\xi t}\sin(\xi x)$ that also adheres to the boundary condition at $x=0$. Choosing $\xi\in\mathbb{Z}$ gives a basis which satisfies the condition at $x=\pi$ as well. Both the calculation of the basis, along the lines of Fourier series, and an implementation are described in Appendix J.1. Even the following theorem holds. ∎
@@
B-EPGP gives $\{e^{\sqrt{-1}jx}\sin(jx)\}_{j\in\mathbb{Z}}$ as basis with dense span in the solutions of (4).
@@
The wave equation in a square with Dirichlet boundary conditions reads:
$\begin{cases}u_{tt}-u_{xx}-u_{yy}=0&\text{for }x,y\in(0,\pi),\,t>0\\ u=0&\text{if $x$ or $y=0$ or $\pi$.}\end{cases}$ (5)
@@
In this case we retrieve the method of separation of variables (a classic Fourier series approach). The calculation of the basis and an implementation are described in Appendix H.1. ∎
@@
B-EPGP gives the basis (6) for (5). The span of (6) is dense in the solutions of (5).
In Theorem D.1 we give a general statement which incorporates both Theorems 3.9 and 3.11.
B-EPGP defines a valid GP and thus offers a fully probabilistic regression model. It not only handles noisy data but also allows sampling from underdetermined PDEs–that is, systems with non-unique solutions despite boundary conditions. In contrast, our previous examples were sufficiently constrained to yield nearly unique solutions. B-EPGP can also be conditioned on arbitrary (noisy or exact) observations. For example, consider the 2D wave equation with Dirichlet boundary conditions:
$\begin{cases}u_{tt}-u_{xx}-u_{yy}=0&\text{for }x,y>0\\ u=0&\text{on }x\in\{0,1\}\text{ and }y\in\{0,1\}\end{cases}$
Figure 1 shows snapshots of a random B-EPGP samples at five timepoints, illustrating its ability to capture uncertainty from sparse data.
@@ 3.3 Hybrid B-EPGP
For domains with both flat and curved pieces of the boundary, we use B-EPGP to fulfill the linear boundary conditions and the direct method from Remark 3.1 using data for the curved pieces.
Example 3.12 (Circular sector).
@@
Consider the space-time domain $\Omega=\{(x,y,t)\colon x^{2}+y^{2}<1,x>0,y>0,t>0\}$ which has the spatial boundary $\Gamma=\left([0,1]\times\{0\}\right)\cup\left(\{0\}\times[0,1]\right)\cup\{(\cos\theta,\sin\theta)\colon\theta\in[0,\pi/2]\}$. We consider the 2D wave equation in this circular sector, $u_{tt}-u_{xx}-u_{yy}=0\text{ in }\Omega$ with Dirichlet boundary condition $u=0$ on $\Gamma$. To account for the flat pieces of $\Gamma$, B-EPGP gives a basis similar to the one in Example 3.7, namely
$e^{\alpha x+\beta y+\tau t}-e^{-\alpha x+\beta y+\tau t}-e^{\alpha x-\beta y+\tau t}+e^{-\alpha x-\beta y+\tau t}\quad\text{for $\alpha,\beta,\tau\in\mathbb{C}$ and $\alpha^{2}+\beta^{2}=\tau^{2}$.}$
@@
To account for the circular piece, we sample many points $(\cos\theta,\sin\theta,t)$ and restrict our GP to satisfy $u=0$ at those points. For experiments, see Section 5.3. ∎
4 Related ML Approaches to PDEs
The intersection of machine learning and PDEs has grown into a vibrant research field with a variety of modeling paradigms. In this section, we survey classes of methods, including neural operators, GPs, and hybrid models, placing our B-EPGP framework in context.
A dominant class of approaches uses neural networks either to approximate the solution directly or as implicit function approximators. A foundational work in this domain is that of physics-informed neural networks (PINNs) (raissi2019physics, ), which penalize deviation from the PDE operator in the loss function. Variants and improvements of this idea abound, such as variational PINNs (Kharazmi2021, ), Fourier PINNs (wang2021eigenvector, ), and gradient-enhanced PINNs (yu2022gradient, ). These approaches train neural networks to minimize the PDE and approximate data at the same time. For a similar approach to GPs, see (chen2022apik, ).
Physics-informed neural networks with hard constraints—often called Ansatz PINNs—enforce boundary and initial conditions exactly by constructing a trial solution $\hat{u}=g+h\,N_{\theta}$, where $g$ satisfies the prescribed conditions and $h$ vanishes on the constraint sets (Lagaris1997, ). Recent advances make this strategy practical on complex geometries via smooth signed/approximate distance functions, $R$-functions, and transfinite interpolation, enabling exact Dirichlet enforcement and robust handling of mixed boundary types (Sukumar2022, ; EPINN2023, ). Relative to penalty-based (“soft”) formulations, hard-constraint designs reduce loss-weight tuning, improve optimization conditioning, and often train faster and more stably, while remaining compatible with data terms and inverse problems (Heliyon2023, ; ExactVsSoft2023, ). Extensions cover Neumann/Robin conditions (via tailored derivative constraints or energetic forms), high-order PDEs (first-order reformulations), and time-dependent problems (time-factorized ansätze), establishing hard-constraint PINNs as a strong default when admissible trial spaces can be constructed (FOPINN2022, ; ATPINNHC2025, ). In contrast to these approaches, B-EPGP exactly satisfies both boundary conditions and the system of PDEs.
Neural operators (kovachki2023neural, ), such as DeepONets (Lu2021deeponet, ) and Fourier Neural Operators (FNOs) (Li2021, ), represent mappings between function spaces and are trained across families of PDEs. Most neural operators face difficulties enforcing hard boundary constraints, especially in high-dimensional or complex geometries. Recent contributions have tackled multi-scale problems (Fang2023solving, ) and high-frequency regimes (raonic2024convolutional, ). Neural operators are connected to GPs (magnani2024linearization, ) via linearization. The approaches (hennig2015probabilistic, ; pfortner2022physics, ; schober2014probabilistic, ; zhang2025monte, ) improve upon numerical PDE-solvers by casting them in a probabilistic framework.
GP models have been applied to both forward and inverse problems involving PDEs. Classical works encode linear constraints through covariance functions (MacedoCastro2008, ; MagneticFieldGP, ; sarkka2011linear, ). Specifically, WIGPR (henderson2021stochastic, ) considers the 3D wave equation. Later methods use symbolic tools to construct priors consistent with linear PDEs (jidling2018probabilistic, ; LH_AlgorithmicLinearlyConstrainedGaussianProcesses, ; LinearlyConstrainedGP, ). Theoretical results underlying these approaches can be found in (henderson2023characterization, ; sullivan2024hille, ) and a related approach to neural networks in (hendriks2020linearly, ; ranftl2023physics, ). Boundary conditions in GPs with differential equations have been modeled via vertical scaling with polynomial (LH_AlgorithmicLinearlyConstrainedGaussianProcessesBoundaryConditions, ) and analytic (langehegermann2022boundary, ) functions. A special case was later introduced as BCGP (dalton2024boundary, ). Well-posed boundary values problems where only observations of the source term exist and no observations of the solution have been considered with GPs using spectral expansions of covariance functions (gulian2022gaussian, ), including Lie symmetries (dalton2024physics, ). EPGP has recently been generalized to inverse problems (li2025gaussian, ).
There is an extensive treatment of Linear PDE coming from numerics and optimization, including finite element methods (larson2013finite, ; logan2011first, ; johnson2009numerical, ; brenner2008mathematical, ; braess2001finite, ), finite difference methods (ervedoza2012wave, ; zuazua2005propagation, ; langtangen2016finite, ; langtangen2017finite, ; thomas2013numerical, ), and both (cohen2017finite, ; mazumder2015numerical, ; ames2014numerical, ; esfandiari2017numerical, ). We do not compare with them since they use deterministic models for well-posed problems, while we construct probabilistic models for under-determined problems.
Physics informed machine learning models play an important role in ODE theory and control. (pmlr-v120-geist20a, ) constructs GP models for control for rigid body dynamics, while (besginow2022constraining, ) constructs such models for general linear ODE systems with constant coefficients, leading to control approaches (besginow2025linear, ; tebbe2024physics, ).
5 Experimental Comparison
B-EPGP is, as a GP, a probabilistic framework for describing the solution set of certain PDE systems. Our Theorems and their proofs show that all realizations of B-EPGP are solutions and moreover that realizations are dense in the set of solutions. To show the usefulness of these theoretical guarantees, we repurposed the B-EPGP priors as solvers by conditioning on enough data. This allows to compare B-EPGP to Neural Operators, the state-of-the-art in solving PDEs from initial values, and EPGP, the state-of-the-art in probabilistic modeling of linear PDEs with constant coefficients. Results for further examples and details can be found in Appendices H.1H.2H.3IJ.1J.3KL.1L.2. These include (i) details on the calculation of the B-EPGP basis and plots for solving the wave and heat equations in several domains in App. HJK, (ii) comparison of our method with EPGP in App. I, (iii) an extension of our method to include inhomogeneous systems $A(\partial)u=f$ and $B(\partial)u=g$ as opposed to our homogeneous case study of (1) in App. L. We measure the accuracy of solutions in two ways: (a) verify the accuracy of our solvers when we know the unique solution to the PDEs, (b) check the conservation of energy, an important physical invariant, for the wave equation in bounded domains (hansen2023learning, ). See Appendix G for details. We refrain from checking the error of a solution in the PDE or boundary condition as it unfairly favors B-EPGP, which satisfies them exactly. These experiments also demonstrate that B-EPGP can be applied in practice under various circumstances.
5.1 Comparison to Neural Operators
Table 1: The median $L_{1}$ error in $0,4]\times[0,8]$ for the experiment from Section 5.1 shows the superiority of B-EPGP for different numbers of data points $n$.
@@ 5.1 Comparison to Neural Operators
BCGP 3.32e-4 6.34e-5 0.62% 0.11%
WIGPR 5.12e-4 8.34e-5 0.86% 0.21%
We compare to the Convolutional Neural Operators (CNO) (raonic2024convolutional, ) and Fourier Neural Operators (FNO) (Li2021, ). As experimental setting we use the 1D wave equation with Neumann boundary condition,
$\begin{cases}u_{tt}=u_{xx}&\text{in $(0,\infty)\times(0,\infty)$}\\ u(0,x)=h_{1}(x)\text{ and }u_{t}(0,x)=h_{2}(x)&\text{for $x\in[0,\infty)$}\\ u_{x}(t,0)=0&\text{for $t\in(0,\infty)$}\end{cases}$
where $h_{1}(x)=f(x-3)+f(x+3)+g(x-1)+g(x+1)$ and $h_{2}(x)=f^{\prime}(x-3)-f^{\prime}(x+3)$ for $f(x)=e^{-5x^{2}}$, $g(x)=e^{-10x^{2}}$. We compare to the unique exact solution given by
$f(x+t-3)+f(x-t+3)+\tfrac{1}{2}\left(g(x+t-1)+g(x-t-1)+g(x+t+1)+g(x-t+1)\right)$
Appendix C encodes the Neumann boundary into B-EPGP, leading to $\{e^{\alpha(x\pm t)}+e^{\alpha(-x\pm t)}\}_{\alpha\in\mathbb{C}}.$ We consider both $n=121$ and $n=1201$ sample points from initial condition on the $t=0$-interval $[0,12]$, and for EPGP, CNO and FNO, we model the boundary at $x=0$ by adding data at $t=0.2\mathbb{Z}_{\geq 0}$ Table 1 shows that EPGP is between one and two orders of magnitude superior to the neural operator methods, and B-EPGP improves upon EPGP, BCG, and WIGPR by factors of around 2. Experimental details about the neural operator computations and ours can be found in Appendix M.1.
5.2 High-dimensional Example: 3D Wave Equation
The 3D wave equation is computationally challenging even without boundary conditions. In Appendix F, we adapt EPGP to fit both initial displacements and velocities. Here, we use B-EPGP with Neumann boundary condition on the planes $y=0$ and $z=0$ and initial conditions as follows.
[…]
where $r_{1},r_{2},r_{3}$ denote distances between $(x,y,z)$ and (1,1,1),(1,1,1),(1,1,1)(1,1,1),(1,-1,1),(1,1,-1), respectively.
The B-EPGP basis can be computed using the ideas in Section J.2 as
$e^{ax+by+cz+dt}+e^{ax-by+cz+dt}+e^{ax+by-cz+dt}+e^{ax-by-cz+dt}$
$+e^{ax+by+cz-dt}+e^{ax-by+cz-dt}+e^{ax+by-cz-dt}+e^{ax-by-cz-dt}\quad\text{for $d^{2}=a^{2}+b^{2}+c^{2}$}.$
Computation of solutions for PDE in 4D incurs a curse of dimensionality, hence few papers tackle 3D wave equation. For instance, EPGP runs out of memory on an Nvidia A100 80GB. Our B-EPGP finishes with a low L1-error of 0.00088. For a comparison of other computation times of machine learning models and finite element solvers, see Appendix A. These machine learning methods take training times in the range of (at least) hours and might even need training data that needs to be generated in the range of days or weeks. However, they have quick inference times (less than a second). FEM solvers are very dependent on the precise circumstances of the setup, but for the above examples usually take minutes (for limited accuracy) to hours. Training B-EPGP (determining the frequencies) takes 2371 seconds, giving a full probabilistic model instead of “just” a single solution. The training details are in Appendix M.2.